OpenStudy (anonymous):

Please help! Equation for a hyperbola with vertices (9,3) and (5,3) and with foci (11,3) and (3,3)?

OpenStudy (anonymous):

a. (y-3)^2/4 - (x-7)^2/12=1 b. (y-3)^2/12 - (y-7)^2/4=1 c. (x-7)^2/12 - (y-3)^2/4=1 d. (x-7)^2/4 - (y-3)^2/12=1

OpenStudy (anonymous):

You ask these questions constantly. When people give you a response it is meant to be a way for You to learn how to complete the problem. Clearly you are just having other people do your homework.

OpenStudy (anonymous):

This is a problem NOT exactly similar to the previous ones I have asked. And I greatly appreciate the help I am given.